I should start a new thread saying WHY DIDN'T I WIN ON EVERY ENTRY?
Comment has been collapsed.
5 Comments - Last post 7 minutes ago by FallenDream
1,531 Comments - Last post 25 minutes ago by simrafael
2,034 Comments - Last post 33 minutes ago by m0r1arty
0 Comments - Created 1 hour ago by malkavian1331
2,256 Comments - Last post 2 hours ago by gudetama1
14 Comments - Last post 3 hours ago by KurganWongraven
20 Comments - Last post 5 hours ago by skycycle
516 Comments - Last post 13 minutes ago by Baph
134 Comments - Last post 21 minutes ago by malachi83
341 Comments - Last post 46 minutes ago by Calibr3
5,997 Comments - Last post 59 minutes ago by igel2005
104 Comments - Last post 1 hour ago by AllTracTurbo
79 Comments - Last post 1 hour ago by DancingXmas
3,044 Comments - Last post 1 hour ago by MouseWithBeer
If you want to cry about entries at least show it in statistically relevant terms. That is, don't tell us you entered 500 giveaways and didn't win, that is irrelevant. If you entered 500 contests with 1/10 odds that would be amazing, more than likely though you entered 500 contests that gave you odds of 1/2500, so that fact that you didn't win makes more sense than if you had.
Here is some easy math so you can go play with yourself in the corner when you realize you have nothing special to say.
Take each contest and divide the number of entrants by the number of gifts, so a giveaway with 1000 copies of dwarfs with 13,000 entrants would be 13. Now divide the number 1 by the number you just solved for. 1/13= .07blahblahblah. This is your odds of winning, less than 8% will win this contest. If you do that for every contest and you add them, that is your cumulative odds, and no that doesn't mean when it equals 1 you should have won, we'll get to that. My cumulative has been over 2.5, so don't bitch.
If you get the inverse of the individual contest odds, 1-1/13= ~92.blahblahblah that represents the odds you will lose, if you multiply every contest since you won last, that would represent the odds that you would have lost every contest, which becomes more relevant to the topic. If you have a total probability of losing that is less than 10% then congrats that would seem relevant over 500 similar events, and still no fucks will be given.
Before any clowns start crap about statistics, this was made simple so people could go waste their time with pointless maths rather than starting even more pointless threads without maths.
*drops pen
Comment has been collapsed.